Sol Summary – May 1st

Crew 193 Sol Summary Report – 1MAY2018

Sol 9

Summary Title: We Sol-emnly Swear We Are Up to No Good, as per the Mars-rauder’s Map

Author: Shawna Pandya, Health and Safety Officer

Mission Status: Nominally nominal

Sol Activity Summary: The day started off later than usual as the EVA-16 team opted to delay their morning excursion until 0900 in the face of threatening storm clouds to the South and East. The morning EVA then progressed as expected, with the PM team keeping up with usual Hab operations, cleaning schedules, GreenHab upkeep and data collection in the interim. Lunch was a Martian feast of curried kidney beans and brown rice (so much better than chianti and fava beans). The EVA-17 team left shortly thereafter to scout the area just west of Zubrin’s head, wisely choosing to turn back 90min ahead of schedule in anticipation of the weather, luckily only meeting with Martian rain for a slight period at the tail end of their sojourn. The crew astronomer managed to capture one solar prominence and a filament. Hab operations otherwise continued as per usual, with one brave crew member opting to try waterless shampoo, and photo-documenting the results (spoiler alert: said crew member still has all her hair).

Look Ahead Plan: EVA 18 will consist of a Green Valley and Mid-Ridge Planitia exploration, whereas EVA 19 will take up an excursion just Southwest of Kissing Camels Ridge. The team continues to function well under the guidance of the dear Commander. Crew development and team-building will continue by way of ASL, team stories, movies and a good ol’ fashioned round of Shanghai, although to be fair, the latter is probably more likely to result in team-breaking than anything else.

Anomalies in work:

1. The crew astronomer was accidentally locked out for a brief spell, but was let in after pleading his case by way of loud noises, and harbors no grudges (we hope).

Weather: Overcast in the AM, rainy and gusty (a Mars-nado, really) in the PM, resulting in a premature return to base on PM EVA.

Crew Physical Status: "So good it hurts," as per one team member.

EVA: EVA-16 proceeded with an exploration of Lith Canyon to the North, and accrued the usual data regarding emergency shelters and crew comfort, safety, health and well-being. Similarly, EVA-17 proceeded with an exploration just West of Zubrin’s Head, and accrued the same data regarding shelters and crew well-being. EVA-17 was halted prematurely due to incrementally inclement weather, which then became overtly inclement.

Reports filed:

GreenHab Report

Astronomy Report

EVA Request

Journalist Report

Science Report

Sol Summary

Photo Report & Photo of the Day

Operations Report

Reports to be filed:

EVA Report

Support Requested:

1. Yogurt. Please send all the yogurt. The crew’s blood levels of Vitamin Y are critically low, and it makes us sad. 🙁

Copyright © The Mars Society. All rights reserved. | Main Site